# The Critical LYZ equation in Kähler Geometry

### v1

@@ 3.2. Key inequalities by the subsolution condition
Let $\mu\in\mathbb{R}^{n}$ satisfy $A(\mu):=\min_{1\leq j\leq n}\sum_{i\neq j}\arctan\mu_{i}>A_{0}>(n-3)\frac{\pi}{2}$. Then for all $\lambda\in\mathbb{R}^{n}$ such that $\theta(\lambda)=\theta(t)$, there exist positive constants $t_{0}$ and $\delta_{0}$ depending only on $n,\mu,A_{0}$ and $C$ in Lemma 2.3 such that for any $t\in(0,t_{0})$, one of the following holds:
@@
    $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}},\ \text{or}$
@@
    $\frac{1}{1+\lambda_{j}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}\ \ \text{for}\ 1\leq j\leq n.$
@@
Since $A(\mu)>A_{0}>(n-3)\frac{\pi}{2}$, we can choose a small positive constant $\delta$ such that
$A(\mu-2\delta{{\bf{1}}})>\frac{A_{0}+(n-3)\frac{\pi}{2}}{2},$
where ${\bf{1}}=(1,\cdots,1)\in\mathbb{R}^{n}$.
For $\lambda\in\mathbb{R}^{n}$ satisfying $\theta(\lambda)=\theta(t)$, we assume $\lambda_{1}\geq\cdots\geq\lambda_{n}$. We know ${\bf{n}}=(\frac{1}{1+\lambda_{1}^{2}},\cdots,\frac{1}{1+\lambda_{n}^{2}})$ is an inward normal vector of $\Gamma^{\theta(t)}$ at $\lambda$. Let $T_{\lambda}$ be the tangent plane of the hypersurface $\partial\Gamma^{\theta(t)}$ at the point $\lambda$.
Case 1: $\big((\mu-\delta{{{\bf 1}}})-\lambda\big)\cdot{\bf{n}}\geq 0$. In this case, we have
@@
Case 2: $\big((\mu-\delta{{{{{\bf 1}}}}})-\lambda\big)\cdot{\bf{n}}<0$. We can show there exists a positive constant $s$ such that $\mu-\delta{{{{{\bf 1}}}}}+se_{1}\in{\Gamma^{\theta(t)}}.$ Indeed, we have
$\theta(\mu-\delta{{{\bf 1}}}+se_{1})=$ $\arctan(\mu_{1}-\delta+s)+\sum_{i=2}^{n}\arctan(\mu_{i}-\delta)$
$>$ $\arctan(\mu_{1}-\delta+s)+A(\mu-\delta{{{\bf 1}}})$
$\geq$ $\arctan(\mu_{1}-\delta+s)+\frac{1}{2}(A_{0}+(n-3)\frac{\pi}{2})$
$>$ $\theta(t)$
if we choose $s=|\delta-\mu_{1}|+\tan\big(\frac{\pi}{2}-\frac{1}{4}(A_{0}-(n-3)\frac{\pi}{2})\big)$ and assume $0<t<t_{0}$ for $t_{0}=\frac{1}{8C}(A_{0}-(n-3)\frac{\pi}{2})$.
The convexity of $\Gamma^{\theta(t)}$ implies $\mu-\delta{{{\bf 1}}}$ and $\mu-\delta{{{\bf 1}}}+se_{1}$ are on two sides of the tangent plane $T_{\lambda}$. Then there exists $a_{0}\in(0,1)$ such that $a_{0}(\mu-\delta{{{\bf 1}}}+se_{1})+(1-a_{0})(\mu-\delta{{{\bf 1}}})=\mu-\delta{{{\bf 1}}}+a_{0}se_{1}\in\mathrm{T_{\lambda}}$ . We also have $\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n}\in\Gamma^{\theta(t)}$ which follows from
$\theta(\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n})$
$=$ $\arctan(\mu_{1}-\delta+a_{0}s+s)$
$+\sum_{i=2}^{n-1}\arctan(\mu_{i}-\delta)+\arctan(\mu_{n}-2\delta)$
$\geq$ $\arctan(\mu_{1}-\delta+a_{0}s+s)+A(\mu-2\delta\bf{1})$
$\geq$ $\arctan(\mu_{1}-\delta+s)+\frac{1}{2}(A_{0}+(n-3)\frac{\pi}{2})$
@@
Thus we have
$0\leq$ $\big((\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n}\big)-(\mu-\delta{{{\bf 1}}}+a_{0}se_{1})\big)\cdot\bf{n}$
$=$ $(se_{1}-\delta e_{n})\cdot\bf n$
@@
or $\frac{1}{1+\lambda_{1}^{2}}\geq\delta s^{-1}\frac{1}{1+\lambda_{n}^{2}}$. Since $\lambda_{1}\geq\cdots\geq\lambda_{n-1}\geq|\lambda_{n}|$ by Lemma 2.2, we get
$\frac{1}{1+\lambda_{j}^{2}}\geq\frac{\delta}{sn}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}},1\leq j\leq n.$ (19)
In view of (18) and (19), if we choose $\delta_{0}=\frac{\delta}{ns}$, we finish the proof. ∎
By the above lemma, we get the following important lemma and the estimate here does not depend on $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
@@
Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6). Then there exist uniform positive constants $t_{0}$ and $\delta_{0}$ such that for any $t\in(0,t_{0})$ and for any $z\in M$, one of the following holds at $z$:
(i) $\sum\limits_{i,j}F^{i\bar{j}}({\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}}})\geq\delta_{0}\sum\limits_{i,j}^{n}F^{i\bar{j}}g_{i\bar{j}},\ \text{or}$
(ii) $\mathcal{F}_{min}\geq\delta_{0}\sum\limits_{i,j}^{n}F^{i\bar{j}}g_{i\bar{j}},$ where $\mathcal{F}_{min}$ is the smallest eigenvalue of the matrix $(F^{i\bar{k}}g_{j\bar{k}})_{n\times n}$.
@@
Let $\mu=(\mu_{1},\cdots,\mu_{n})$ be the eigenvalues of $\chi_{\underline{u}}$ with respect to $\omega$ with $\mu_{1}\geq\cdots\geq\mu_{n}$. For any $z\in M$, we choose the normal coordinates such that at $z$
$g_{i\bar{j}}=\delta_{ij},\ \chi_{t,u^{t}}=\sum\limits_{i=1}^{n}\lambda_{i}\sqrt{-1}dz^{i}\wedge d\bar{z}^{i}\ \text{with}\ \lambda_{1}\geq\cdots\geq\lambda_{n},$
where $\chi_{t,u^{t}}=\chi+t\omega+\sqrt{-1}\partial\bar{\partial}{u^{t}}$. Then at $z$, we have
$F^{i\bar{j}}=\frac{\delta_{i\bar{j}}}{1+\lambda_{i}^{2}}.$ (20)
Let $\delta_{0}$ and $t_{0}$ be the uniform constants in Lemma 3.2. We use Lemma 3.2 to get the proof.
Case 1: $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})\geq\delta_{0}\sum\limits_{i}F^{i\bar{i}}$. Since $F^{1\bar{1}}\leq\cdots\leq F^{n\bar{n}}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$, we use the Schur-Horn theorem [16] to get
$\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})=$ $\sum_{i,j}F^{i\bar{j}}((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}})+t\sum_{i}F^{i\bar{i}}$
$\geq$ $F^{i\bar{j}}((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}})$
$\geq$ $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})$
$\geq$ $\delta_{0}\sum\limits_{i}F^{i\bar{i}}.$
Case 2: $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})<\delta_{0}\sum\limits_{i}F^{i\bar{i}}$. By Lemma 3.2, we get $\mathcal{F}_{min}=\frac{1}{1+\lambda_{1}^{2}}\geq\delta_{0}\sum\limits_{i}\frac{1}{1+\lambda_{i}^{2}}=\delta_{0}\sum\limits_{i,j}F^{i\bar{j}}g_{i\bar{j}}$. ∎
@@ 3.3. The complex Hessian estimate of Hou-Ma-Wu type
We use the auxiliary function introduced by Hou-Ma-Wu [17] who established complex Hessian estimates for the $k$-Hessian equation on compact Kähler manifold to derive the complex Hessian estimates for families of LYZ equations. Here we use the term $w_{i\bar{j}}\xi^{i}\bar{\xi}^{j}$ instead of the largest eigenvalue of $(g^{j\bar{k}}w_{i\bar{k}})_{n\times n}$ and thus the proof seems much more simple. The complex Hessian estimates here do not depend on $({\theta(t)-(n-2)\frac{\pi}{2}})^{-1}$.
@@
Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$. Then there exists uniform positive constants $t_{0}$ and $C$ such that for any $t\in(0,t_{0})$, we have
$\sup_{M}{|\sqrt{-1}\partial\bar{\partial}u^{t}|_{\omega}}\leq C(1+\sup_{M}|\nabla u^{t}|_{\omega}^{2}).$ (21)
@@
Without loss of generality, we assume $\underline{u}=0$. Let $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$. For simplicity we use $u$ instead of $u^{t}$ in the following proof. Denote $S(T^{1,0}M):=\bigcup_{z\in M}\{\xi\in T^{1,0}_{z}M\,|\,|\xi|_{\omega}=1\}$. We consider the auxiliary function on $S(T^{1,0}M)$
$H(z,\xi)=\log(w_{k\bar{l}}\xi^{k}\bar{\xi}^{l})+\varphi(|\nabla u|^{2}_{g})+\psi(u),$
where $\varphi$ and $\psi$ are given by
$\begin{split}\varphi\left(s\right)&=-\frac{1}{2}\log\left({1-\frac{s}{{2K}}}\right),\quad 0\leq s\leq K-1,\\ \psi\left(s\right)&=-A\log\left({1+\frac{s}{{2L}}}\right),\quad-L+1\leq s\leq 0,\end{split}$
with $K=\mathop{\sup}\limits_{M}|\nabla u|_{g}^{2}+1,\ L=\mathop{\sup}\limits_{M}\left|u\right|+1,\ A=2L(C_{0}+1)$ where $C_{0}$ is a uniform positive constant to be determined later. We have
$(4K)^{-1}$ $\leq\varphi^{\prime}\leq(2K)^{-1},\ \varphi^{\prime\prime}=2\big(\varphi^{\prime})^{2},$ (22)
$\frac{1}{2}{A}{L^{-1}}$ $\leq-\psi^{\prime}\leq{A}{L^{-1}},\ \psi^{\prime\prime}=A^{-1}(\psi^{\prime})^{2}.$ (23)
Suppose the function $H$ attains its maximum at $z_{0}$ along the direction $\xi_{0}$. We choose holomorphic coordinates near $z_{0}$ such that
$g_{i\bar{j}}(z_{0})=\delta_{i\bar{j}},\ \partial_{k}g_{i\bar{j}}(z_{0})=0,\text{and}$ (24)
$w_{i\bar{j}}(z_{0})=\lambda_{i}\delta_{i\bar{j}}\ \textup{with}\ \lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$
which forces $\xi_{0}=\frac{\partial}{\partial z_{1}}$. We extend $\xi_{0}$ near $z_{0}$ as $\tilde{\xi}_{0}(z)=(g_{1\bar{1}})^{-\frac{1}{2}}\frac{\partial}{\partial z_{1}}$. Then the function $Q(z)=H(z,\tilde{\xi}_{0}(z))$ attains its local maximum at $z_{0}$. By the maximum principle, we have at $z_{0}$
@@
$0\geq\mathcal{L}(Q)=F^{i\bar{i}}Q_{i\bar{i}}=$ $\lambda_{1}^{-1}F^{i\bar{i}}w_{1\bar{1},i\bar{i}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}$
$+\varphi^{\prime\prime}F^{i\bar{i}}\big|(|\nabla u|^{2})_{i}\big|^{2}+\varphi^{\prime}F^{i\bar{i}}\big(|\nabla u|^{2}\big)_{i\bar{i}}$
$+\psi^{\prime\prime}F^{i\bar{i}}|u_{i}|^{2}+\psi^{\prime}F^{i\bar{i}}u_{i\bar{i}}.$ (26)
We will estimate each term in (3.3). In the following, we always assume
@@
otherwise we finish the proof.
Since $d\chi=0$, by covariant derivative formulae, we have
@@
By (17) and $\lambda_{i}+\lambda_{j}\geq 0$ for $i\neq j$ by Lemma 2.2, we get
$F^{i\bar{i}}w_{i\bar{i},1\bar{1}}=$ $\sum_{i,j}\frac{\lambda_{i}+\lambda_{j}}{(1+\lambda_{i}^{2})(1+\lambda_{j}^{2})}|w_{i\bar{j},1}|^{2}$

### v2

@@ 3.2. Key inequalities from the subsolution condition
Let $\mu\in\mathbb{R}^{n}$ satisfy
$A(\mu):=\min_{1\leq j\leq n}\sum_{i\neq j}\arctan\mu_{i}\;>\;A_{0}\;>\;(n-3)\frac{\pi}{2}.$
Then, for every $\lambda\in\mathbb{R}^{n}$ with $\theta(\lambda)=\theta(t)$, there exist positive constants $t_{0}$ and $\delta_{0}$, depending only on $n$, $\mu$, $A_{0}$, and the constant $C$ from Lemma 2.3, such that for any $t\in(0,t_{0})$, one of the following statements holds:
@@
    $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}$; or
@@
    $\frac{1}{1+\lambda_{j}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}$ for $1\leq j\leq n$.
@@
Given that $A(\mu)>A_{0}>(n-3)\frac{\pi}{2}$, we can choose a small positive constant $\delta>0$ such that
$A(\mu-2\delta\mathbf{1})>\frac{A_{0}+(n-3)\frac{\pi}{2}}{2},$
where $\mathbf{1}=(1,\ldots,1)\in\mathbb{R}^{n}$.
For $\lambda\in\mathbb{R}^{n}$ satisfying $\theta(\lambda)=\theta(t)$, assume without loss of generality that $\lambda_{1}\geq\cdots\geq\lambda_{n}$. Recall that the vector
$\mathbf{n}=\Bigl(\frac{1}{1+\lambda_{1}^{2}},\cdots,\frac{1}{1+\lambda_{n}^{2}}\Bigr)$
is an inward normal to $\Gamma^{\theta(t)}$ at $\lambda$. Let $T_{\lambda}$ denote the tangent plane to the hypersurface $\partial\Gamma^{\theta(t)}$ at $\lambda$.
Case 1: $\bigl((\mu-\delta\mathbf{1})-\lambda\bigr)\cdot\mathbf{n}\geq 0$. In this case, we have
@@
Case 2: $\bigl((\mu-\delta\mathbf{1})-\lambda\bigr)\cdot\mathbf{n}<0$. We claim that there exists a positive constant $s$ such that $\mu-\delta\mathbf{1}+se_{1}\in\Gamma^{\theta(t)}$. Indeed, observe that
$\theta(\mu-\delta\mathbf{1}+se_{1})$ $=\arctan(\mu_{1}-\delta+s)+\sum_{i=2}^{n}\arctan(\mu_{i}-\delta)$
$>\arctan(\mu_{1}-\delta+s)+A(\mu-\delta\mathbf{1})$
$\geq\arctan(\mu_{1}-\delta+s)+\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr).$
To ensure $\theta(\mu-\delta\mathbf{1}+se_{1})>\theta(t)$, it suffices to choose $s$ satisfying
$\arctan(\mu_{1}-\delta+s)>\theta(t)-\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr).$
Take
$s=|\delta-\mu_{1}|+\tan\Bigl(\frac{\pi}{2}-\frac{1}{4}\bigl(A_{0}-(n-3)\frac{\pi}{2}\bigr)\Bigr),$
and assume $0<t<t_{0}$ with $t_{0}=\frac{1}{8C}\bigl(A_{0}-(n-3)\frac{\pi}{2}\bigr)$. Then the desired inequality holds.
The convexity of $\Gamma^{\theta(t)}$ implies that the points $\mu-\delta\mathbf{1}$ and $\mu-\delta\mathbf{1}+s\mathbf{e}_{1}$ lie on opposite sides of the tangent plane $T_{\lambda}$. Hence, there exists $a_{0}\in(0,1)$ such that the convex combination
$a_{0}(\mu-\delta\mathbf{1}+s\mathbf{e}_{1})+(1-a_{0})(\mu-\delta\mathbf{1})=\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}$
lies on $T_{\lambda}$.
We now show that $\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n}\in\Gamma^{\theta(t)}$. Indeed,
$\theta\bigl(\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n}\bigr)$
$=\arctan(\mu_{1}-\delta+a_{0}s+s)+\sum_{i=2}^{n-1}\arctan(\mu_{i}-\delta)+\arctan(\mu_{n}-2\delta)$
$\geq\arctan(\mu_{1}-\delta+a_{0}s+s)+A(\mu-2\delta\mathbf{1})$
$\geq\arctan(\mu_{1}-\delta+s)+\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr)$
@@
Thus, we have
$0$ $\leq\bigl((\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n})-(\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1})\bigr)\cdot\mathbf{n}$
$=(s\mathbf{e}_{1}-\delta\mathbf{e}_{n})\cdot\mathbf{n}$
@@
which is equivalent to
$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\cdot\frac{1}{1+\lambda_{n}^{2}}.$ (19)
By Lemma 2.2, we have $\lambda_{1}\geq\cdots\geq\lambda_{n-1}\geq|\lambda_{n}|$. Consequently, for each $j=1,\dots,n$,
$\frac{1}{1+\lambda_{n}^{2}}\geq\frac{1}{1+\lambda_{j}^{2}}.$
Combining with (19), we obtain
$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\cdot\frac{1}{1+\lambda_{j}^{2}},\quad\textup{for all $1\leq j\leq n$.}$
Summing over $j=1,\dots,n$ yields
$n\cdot\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\sum_{j=1}^{n}\frac{1}{1+\lambda_{j}^{2}},$
and therefore
$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{ns}\sum_{j=1}^{n}\frac{1}{1+\lambda_{j}^{2}}.$
Since $\lambda_{1}^{2}\geq\lambda_{j}^{2}$ for all $j$, we have
$\frac{1}{1+\lambda_{j}^{2}}\geq\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{ns}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}.$ (20)
In view of (18) and (20), choosing $\delta_{0}=\frac{\delta}{ns}$ completes the proof. ∎
By the preceding lemma, we obtain the following key result, noting that the estimate does not depend on $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
@@
Let $\underline{u}$ be a subsolution of the critical LYZ equation (4) and $u^{t}$ the solution of the LYZ equation (6). Then there exist uniform positive constants $t_{0}$ and $\delta_{0}$ such that for any $t\in(0,t_{0})$ and any $z\in M$, one of the following holds at $z$:
  1. (i)
    $\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})\geq\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}$, or
  2. (ii)
    $\mathcal{F}_{\min}\geq\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}$, where $\mathcal{F}_{\min}$ denotes the smallest eigenvalue of the Hermitian matrix $\bigl(F^{i\bar{k}}g_{j\bar{k}}\bigr)_{n\times n}$.
@@
Let $\mu=(\mu_{1},\dots,\mu_{n})$ be the eigenvalues of $\chi_{\underline{u}}$ with respect to $\omega$, arranged so that $\mu_{1}\geq\cdots\geq\mu_{n}$.
Fix $z\in M$. We choose normal coordinates at $z$ such that
$g_{i\bar{j}}=\delta_{ij},\quad\chi_{t,u^{t}}=\sum_{i=1}^{n}\lambda_{i}\sqrt{-1}\,dz^{i}\wedge d\bar{z}^{i}\quad\text{with}\quad\lambda_{1}\geq\cdots\geq\lambda_{n},$
where $\chi_{t,u^{t}}=\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t}$. In these coordinates, at $z$ we have
$F^{i\bar{j}}=\frac{\delta_{ij}}{1+\lambda_{i}^{2}}.$ (21)
Let $\delta_{0}$ and $t_{0}$ be the uniform constants given by Lemma 3.2. We now prove the lemma by distinguishing two cases.
Case 1: $\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})\geq\delta_{0}\sum_{i}F^{i\bar{i}}$.
Since $F^{1\bar{1}}\leq\cdots\leq F^{n\bar{n}}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$, the Schur–Horn theorem [17] implies
$\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})$ $=\sum_{i,j}F^{i\bar{j}}\bigl((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}}\bigr)+t\sum_{i}F^{i\bar{i}}$
$\geq\sum_{i,j}F^{i\bar{j}}\bigl((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}}\bigr)$
$=\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})$
$\geq\delta_{0}\sum_{i}F^{i\bar{i}}.$
Case 2: $\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})<\delta_{0}\sum_{i}F^{i\bar{i}}$.
By Lemma 3.3, we obtain
$\mathcal{F}_{\min}=\frac{1}{1+\lambda_{1}^{2}}\geq\delta_{0}\sum_{i}\frac{1}{1+\lambda_{i}^{2}}=\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}.$
This completes the proof. ∎
@@ 3.3. The complex Hessian estimate of Hou-Ma-Wu type
We use the auxiliary function introduced by Hou-Ma-Wu [18], who established complex Hessian estimates for the $k$-Hessian equation on compact Kähler manifolds, to derive the complex Hessian estimates for families of LYZ equations. Here, instead of using the largest eigenvalue of $(g^{j\bar{k}}w_{i\bar{k}})_{n\times n}$, we employ the term $w_{i\bar{j}}\xi^{i}\bar{\xi}^{j}$, which simplifies the proof considerably.
@@
Let $\underline{u}$ be a subsolution of the critical LYZ equation (4) and let $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$. Then there exist uniform positive constants $t_{0}$ and $C$, independent of $\bigl(\theta(t)-(n-2)\frac{\pi}{2}\bigr)^{-1}$, such that for any $t\in(0,t_{0})$,
$\sup_{M}|\sqrt{-1}\partial\bar{\partial}u^{t}|_{\omega}\leq C\bigl(1+\sup_{M}|\nabla u^{t}|_{\omega}^{2}\bigr).$ (22)
@@
Without loss of generality, we assume $\underline{u}=0$. Let $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$. For simplicity, we write $u$ instead of $u^{t}$ in the following proof. Denote $S(T^{1,0}M):=\bigcup_{z\in M}\{\xi\in T^{1,0}_{z}M\mid|\xi|_{\omega}=1\}$. We consider the auxiliary function on $S(T^{1,0}M)$:
$H(z,\xi)=\log\bigl(w_{k\bar{l}}\xi^{k}\bar{\xi}^{l}\bigr)+\varphi\bigl(|\nabla u|_{g}^{2}\bigr)+\psi(u),$
where $\varphi$ and $\psi$ are defined by
$\varphi(s)$ $=-\frac{1}{2}\log\!\left(1-\frac{s}{2K}\right),\quad 0\leq s\leq K-1,$
$\psi(s)$ $=-A\log\!\left(1+\frac{s}{2L}\right),\quad-L+1\leq s\leq 0,$
with
$K=\sup_{M}|\nabla u|_{g}^{2}+1,\quad L=\sup_{M}|u|+1,\quad A=2L(C_{0}+1),$
and $C_{0}$ is a uniform positive constant to be determined later. We have the following derivatives:
$(4K)^{-1}$ $\leq\varphi^{\prime}\leq(2K)^{-1},\quad\varphi^{\prime\prime}=2(\varphi^{\prime})^{2},$ (23)
$\frac{1}{2}AL^{-1}$ $\leq-\psi^{\prime}\leq AL^{-1},\quad\psi^{\prime\prime}=A^{-1}(\psi^{\prime})^{2}.$ (24)
Suppose $H$ attains its maximum at $(z_{0},\xi_{0})$. We choose holomorphic coordinates near $z_{0}$ such that
$g_{i\bar{j}}(z_{0})=\delta_{ij},\quad\partial_{k}g_{i\bar{j}}(z_{0})=0,\quad\text{and}$
$w_{i\bar{j}}(z_{0})=\lambda_{i}\delta_{ij}\quad\text{with}\quad\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n},$
which forces $\xi_{0}=\frac{\partial}{\partial z_{1}}$. Extend $\xi_{0}$ near $z_{0}$ by $\tilde{\xi}_{0}(z)=(g_{1\bar{1}})^{-\frac{1}{2}}\frac{\partial}{\partial z_{1}}$. Then the function
$Q(z)=H(z,\tilde{\xi}_{0}(z))$
attains a local maximum at $z_{0}$.
By the maximum principle, at $z_{0}$ we have
@@
$0\geq\mathcal{L}(Q)=\sum_{i}F^{i\bar{i}}Q_{i\bar{i}}=$ $\sum_{i}\lambda_{1}^{-1}F^{i\bar{i}}w_{1\bar{1},i\bar{i}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}$
$+\sum_{i}\varphi^{\prime\prime}F^{i\bar{i}}\big|(|\nabla u|^{2})_{i}\big|^{2}+\varphi^{\prime}\sum_{i}F^{i\bar{i}}\big(|\nabla u|^{2}\big)_{i\bar{i}}$
$+\psi^{\prime\prime}\sum_{i}F^{i\bar{i}}|u_{i}|^{2}+\psi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}.$ (26)
We will estimate each term in (3.3). In what follows, we always assume
@@
otherwise the desired estimate holds trivially.
Since $d\chi=0$, the commutation formula for covariant derivatives yields
@@
By (17) and the fact that $\lambda_{i}+\lambda_{j}\geq 0$ for $i\neq j$ (Lemma 2.2), we obtain
$\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}$ $=\sum_{i,j}\frac{\lambda_{i}+\lambda_{j}}{(1+\lambda_{i}^{2})(1+\lambda_{j}^{2})}|w_{i\bar{j},1}|^{2}$